OpenStudy (anonymous):

how long will it take for an investment of $1,000 to double in value if the interest rate is 8.5% per year compounded continuously?

ganeshie8 (ganeshie8):

2000 = 1000*e^(.085*t) can you solve for t by taking log both sides

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I got a different answer so I messed up somewhere...

ganeshie8 (ganeshie8):

2 = e^(.085t) ln2 = .085t t = ln2/.085 = ~ 8.15 years

ganeshie8 (ganeshie8):

hmm.. hw much u got

OpenStudy (anonymous):

How is the log of 2000 2?

ganeshie8 (ganeshie8):

2000 = 1000*e^(.085*t) dividing both sides by 1000 2 = e^(.085t)

OpenStudy (anonymous):

k i got it thanks!

ganeshie8 (ganeshie8):

brilliant ! yw ^_^